OpenStudy (anonymous):

find the exact value of 10pi/4

OpenStudy (anonymous):

\[\frac{ 10\pi }{ 4 }\]

OpenStudy (anonymous):

7.85398163397

OpenStudy (anonymous):

that's not one of my choices

OpenStudy (anonymous):

hmm im not sure and i dont want give you the wrong answer

OpenStudy (anonymous):

my mistake, it's \[ \sin \frac{ 10\pi }{ 4 }\]

OpenStudy (anonymous):

since you are given decimal approximations as answers, you can use a calculator

OpenStudy (anonymous):

when i use my calculator i get .136

OpenStudy (anonymous):

or you can say that \(\frac{10\pi}{4}=\frac{5\pi}{2}\) which is coterminal with \(\frac{\pi}{2}\) and \[\sin(\frac{\pi}{2})=1\]

OpenStudy (anonymous):

make sure your calculator is in "radian" mode then you will get 1

OpenStudy (anonymous):

okay thanks a ton!